OpenStudy (anonymous):

i learned this last year, but i forgot..how do you find percent..for example 135 is "x" percent of 265?

jimthompson5910 (jim_thompson5910):

135 is "x" percent of 265 translates to 135 = (x/100)*265

jimthompson5910 (jim_thompson5910):

now solve for x 135 = (x/100)*265 100*135 = 265x 13500 = 265x 265x = 13500 keep going to solve for x

OpenStudy (whpalmer4):

In effect, divide 135/265 and move the decimal point 2 places to the right. For example, if you got 4 out of 10 right, that would be 4/10 = 0.40 or 40 %.

OpenStudy (whpalmer4):

It's interesting that another kid is also asking how to figure out their grade, allegedly having gotten 135 points out of 265 :-)